# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v3

@@ 4.3 Three-phase estimation procedure
    Estimate the autoregressive and weight function parameters $\boldsymbol{\theta}_{1}\equiv(\phi_{1,0},...,\phi_{M,0},\varphi_{1},...,\varphi_{M},\alpha)$ by the method of (penalized) nonlinear least squares (NLS) as described in Appendix C. Denote this estimate of $\boldsymbol{\theta}_{1}$ as $\hat{\boldsymbol{\theta}}_{1}^{NLS}$.
@@
Step 1 obtains initial estimates for the AR and weight function parameters, and it is comparable to Step 1 of Lanne et al. (2017) but with the nonlinearity of the model explicitly accounted for. Notably, the NLS estimation often produces estimates that do not satisfy the usual stability condition for each of the regimes, and thus allowing for instability (but penalizing it) facilitates utilization of the NLS estimates in our three-phase procedure. Step 2 is comparable to Step 2 of Lanne et al. (2017), but the multimodality of the (penalized) log-likelihood function is taken into account by making use of a robust estimation algorithm that is able escape from local maxima. Step 3 finalizes the estimation by initializing a standard gradient based optimization algorithm from the initial estimates obtained from the previous steps, often converging to a nearby local maximum of the (penalized) log-likelihood function.
@@
Due to the high multimodality of the (penalized) log-likelihood function, Steps 2 and 3 should be ran a large number times to improve the reliability of the results. Since multiple estimation rounds are ran, the estimation procedure produces a set of estimates, possibly corresponding to different local maximums of the (penalized) log-likelihood function. If there are multiple solutions with (penalized) log-likelihoods almost equal to the highest found (penalized) log-likelihood, the blended identification procedure discussed in Section 3.5 can be employed to obtain the best local solution that facilitates labelling the shocks of interest. The proposed three-phase estimation procedure is implemented to the accompanying R package sstvars (Virolainen, 2025a), which is available through the CRAN repository.
@@
To assess the performance of the penalized maximum likelihood (PML) estimator as well as our three-phase estimation procedure, we conduct a small-scale Monte Carlo study, which is discussed in detail in Appendix D. Since estimation is computationally demanding, we consider simple bivariate (structural) TVAR and LSTVAR models with $p=1$ and $M=2$. To evaluate how the PML estimator performs when the AR matrices are in the penalization region, we consider two specifications of both TVAR and LSTVAR models: one where the AR matrices are well inside the stability region and one where they lie clearly in the region where the penalization term is strictly positive.
@@
According to the results (Tables D.2 and D.3 in Appendix D), estimation accuracy is slightly better when the AR matrices are inside the stability region, but even when they are in the penalization region, the estimators of the AR matrices exhibit only minor small-sample bias. Additionally, estimation accuracy appears to be slightly higher for the TVAR models compared to the LSTVAR models. While for most of the parameters there is only very slight bias, we find a notable upward small-sample bias for the scale parameter, for a degrees-of-freedom parameter, and for a certain entry in each impact matrix, explainable by the statistical properties of the true data generating process. However, the bias diminishes for all of the parameters as the sample size increases. Since also the standard deviations of the estimates diminish as the sample size increases, our results are in line with consistency of the PML estimator. Overall, the performance of the proposed PML estimator and the three-phase estimation procedure seems reasonable.
@@ 5 Empirical application
Our empirical application studies the macroeconomic effects of the climate policy uncertainty (CPU) shock. In the related literature, Fried et al. (2022) develop a dynamic general equilibrium model incorporating beliefs about future climate policy, and find that an increased climate policy transition risk shifts investments toward cleaner capital while reducing total investments and output. Khalil and Strobel (2023) reach similar conclusions in their DSGE model, attributing the effects to financial institutions’ aversion to uncertainty. Huang and Punzi (2024) also use a DSGE model and find that CPU shocks cause firms to delay investments, reducing output and increasing inflation, which is supported by results from their recursively identified Bayesian VAR.
We are interested in studying how the effects of the CPU shock vary depending on the level of economic policy uncertainty (EPU). Such variation can be accommodated in our structural STVAR model by specifying a logistic transition weight function with the (lagged) level of EPU as the switching variable. This approach allows us to capture potential nonlinearities and regime-dependent effects, providing more detailed results on the macroeconomic effects of the CPU shock than linear SVAR analysis.
@@ 5.1 Data
As a measure of real economic activity, we include the log of industrial production index (IPI), which is detrended by taking first differences. For measuring the price level, we use the logs of consumer price index (CPI) and producer price index (PPI), likewise detrended by taking first differences. Finally, we measure the monetary policy stance with the effective Federal funds rate that is replaced by the Wu and Xia (2016) shadow rate for the zero-lower-bound periods (RATE). The series of the included variables are presented in Figure 1.
Figure 1: Monthly U.S. time series covering the period from 1987:4 to 2024:12. From top to bottom, the variables are the climate policy uncertainty index (divided by $10$); economic policy uncertainty index (divided by $10$); the log-differences of industrial production index, consumer price index, and producer price index (multiplied by $100$); and the effective federal funds rate replaced by the Wu and Xia (2016) shadow rate for the zero lower bound periods. The bottom panel shows the transition weights of High EPU Regime (Regime 2) of the fitted two-regime second-order structural LSTVAR model. The shaded areas indicate the U.S. recessions defined by the NBER.
@@ 5.2 The structural LSTVAR model
In order to specify the structural STVAR model (3.2)-(3.3), the transition weights and the distributions of the structural shocks must be specified. We employ the logistic transition weights with two regimes (Anderson and Vahid, 1998) given in Equation (2.4). They facilitate smooth transitions between the regimes depending the level of the switching variable, enabling an analysis of how the effects of the CPU shock vary depending on the level of EPU by specifying its first lag as the switching variable. The logistic weight function also has the convenient property that is nests the threshold weight function (Tsay, 1998) as a special case when the scale parameter tends to infinity, effectively incorporating selection between these two weight functions within estimation.
@@
We estimate the structural LSTVAR model by PML with the three-phase procedure proposed in Section 4, assuming that the structural shocks follow the independent skewed $t$-distributions. The skewed $t$-distribution can flexibly capture fat tails and skewness, and as the Gaussian distribution is obtained as a special case of the skewed $t$-distribution, with $\nu_{i}=\infty$ and $\lambda_{i}=0$, the validity of the identifying Assumption 1 can be assessed based on the estimates of these parameters. To avoid overfitting, particularly in the regime prevailing when the level of EPU is high, the autoregressive order $p=2$ is selected based on Hannan-Quinn information criterion.
@@
To deal with the problem of weak identification with respect to the ordering and signs of the columns of $B_{2}$, discussed in Section 3.3, 3.4, and 3.5, we implement the following, simplistic blended identification strategy that avoids making restrictive assumptions about the effects of the CPU shock. We normalize the first row of $B_{1}$ positive with CPUI ordered first and assume that in both regimes, one of the shocks, deemed as the CPU shock, has greater impact effect on the CPUI than any other shock, and this impact effect has the same sign in both regimes. Then, to support our identification, we check that the recovered CPU shocks align with several major historical events affecting CPU. In particular, we find that President Trump’s election is preceded by large CPU shocks in October 2016 and 2024 ($17$ and $26$, respectively). The largest negative recovered CPU shock ($-10$), in turn, dates to August 2022, when the Inflation Reduction Act was passed, providing support for the clean energy transition. Among all the shocks recovered from the model, only the one deemed as the CPU shock has signs consistent with the above historical events affecting CPU.
@@
The location parameter estimate is $\hat{c}=14.48$, implying roughly that Regime 1 dominates when the level of EPU is not high, while Regime 2 starts prevailing in times of high EPU. Therefore, we label Regime 1 as Low EPU Regime and Regime 2 as High EPU Regime. The scale parameter estimate is $\hat{\gamma}=16.65$, indicating that transitions between the regimes are somewhat fast (rather than very smooth), which can be observed also from the bottom panel of Figure 1 where the fitted transition weights of High EPU Regime are presented.
@@
The estimates of the degrees-of-freedom and skewness parameters are (2.33,2.10,2.64,3.23,6.54,(2.33,2.10,2.64,3.23,6.54,( 2.33 , 2.10 , 2.64 , 3.23 , 6.54 , 2.68)2.68)2.68 ) and $(0.41,0.17,0.09,0.43,0.25,-0.12)$, respectively. Clearly, none of the shocks is close to Gaussian, suggesting that Assumption 1 is satisfied. Then, we check whether our model satisfies the stationarity condition of Theorem A.1 in Appendix A by computing an upper bound for the joint spectral radius of the companion form AR matrices of the regimes using the JSR toolbox (Jungers, 2023) in MATLAB. The obtained upper bounds is $0.98$, which is strictly less than one, implying that our model satisfies Assumption A.1 (in Appendix A) and is ergodic stationary by Theorem A.1. Based on the residual diagnostics presented in Appendix E, the adequacy of our model to capture the autocorrelation structure, conditional heteroskedasticity, and distribution of the data seems, overall, reasonable.
@@ 5.3 Generalized impulse response functions
Due to the $p$-step Markov property of our structural LSTVAR model, conditioning on the previous $p$ observations, $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$, is equivalent to conditioning on $\mathcal{F}_{t-1}$. To the study the state-dependence of the effects of the CPU shock, we follow a procedure similar to Lanne and Virolainen (2025) and calculate the GIRFs conditional on a given regime dominating when the shock arrives. Specifically, for each Regime $\tilde{m}\in\{1,2\}$, we take all the length $p$ histories $\boldsymbol{y}_{t-1}$ from the data for which the corresponding transition weight $\alpha_{\tilde{m},t}$ is greater than $0.75$, indicating that Regime $\tilde{m}$ is clearly dominant (there are $366$ such histories for Low EPU Regime and $80$ for High EPU Regime). For each such history, we take the corresponding CPU shock recovered from the data, and compute the GIRF using the Monte Carlo algorithm described in Lanne and Virolainen (2025), Appendix B. Finally, GIRFs corresponding to shocks of different sign and size are made comparable by scaling them to correspond to a $5$ point instantaneous increase of CPUI.
@@
Figure 2: Generalized impulse response functions to the identified CPU shock $h=0,1,....,36$ months ahead. From top to bottom, the responses of CPUI, EPUI, IPI, CPI, PPI, the interest rate variable, and transition weights of High EPU Regime are depicted in each row. The responses of IPI, CPI, and PPI are accumulated to (100×100$\times$100 ×)log-levels. The left column shows the responses to conditionally on High EPU Regime dominating when the shock arrives, whereas the right column shows the responses conditionally on High EPU Regime dominating. All the GIRFs have been scaled so that the instantaneous increase in CPUI is $5$ points. Each GIRF is computed with a Monte Carlo algorithm similar to the one described in Lanne and Virolainen (2025, Appendix B).
@@
The distribution of the GIRFs in each regime is illustrate in a so-called ”shotgun plot” in Figure 2, which depicts each GIRFs using a level of opacity such that the darkness of a region displays the concentration of GIRFs in it (cf. Lanne and Virolainen, 2025). The response of CPUI to a positive CPU shock follows a similar pattern in both regimes in the vast majority of the GIRFs, while EPUI temporarily increases in Low EPU Regime and decreases in High EPU Regime. A possible explanation for the decrease in the EPUI could be that under already heightened EPU, an unexpected increase in CPU may induce a shift in focus from broader economic policy issues to climate policy.
@@
Consistent with the previous literature (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024), a positive CPU shock decreases production in both regimes. Huang and Punzi (2024) attribute the decrease in production to firms postponing investments, whereas Fried et al. (2022) and Khalil and Strobel (2023) consider mechanisms that shift investments toward cleaner capital. Although many of the GIRFs exhibit a short-term increase of IPI in High EPU Regime, the negative long-term response of IPI seems, on average, roughly similar in both of the regimes.
@@
In both regimes, a positive CPU shock increases consumer prices, aligning with Huang and Punzi (2024) and Khalil and Strobel (2023). Huang and Punzi (2024) found that a positive CPU shock increases inflation by acting as a positive supply shock, whereas according Khalil and Strobel (2023), a positive CPU shock may decrease aggregate demand due to risk-averse households raising savings, while firms raise their prices pre-emptively to lower the probability of being stuck with a negative markup. However, we find that the inflationary effects are stronger in High EPU Regime than in Low EPU Regime. One possible explanation is that the firms are more risk averse in the periods of high EPU, thereby reacting more sensitively and passing the increased costs or markups onto consumers. Producer prices, in turn, mostly decrease in Low EPU Regime but increase in High EPU Regime.
@@
The response of the interest rate variable seems consistent with the inflationary effects of the CPU shock: it increases both regimes, but clearly more so in High EPU Regime. Finally, consistent with the GIRFs of EPUI, a positive CPU shock temporarily increases the transition weights of High EPU Regime in the periods of low EPU, while the effect is opposite in the periods of high EPU. Overall, our results suggest that a positive CPU shock decreases production and increases inflation in both regimes, but its inflationary effects are stronger in the periods of high EPU than in times of low EPU.

### v4

@@ 4.3 Three-step estimation procedure
    Estimate the autoregressive and weight function parameters $\boldsymbol{\theta}_{1}\equiv(\phi_{1},...,\phi_{M},\varphi_{1},...,\varphi_{M},\alpha)$ by the method of (penalized) nonlinear least squares (NLS) (see Appendix C for details). Denote this estimate of $\boldsymbol{\theta}_{1}$ as $\hat{\boldsymbol{\theta}}_{1}^{NLS}$.
@@
Step 1 obtains initial estimates for the AR and weight function parameters, and it is comparable to Step 1 of Lanne et al. (2017) but with the nonlinearity of the model explicitly accounted for. Notably, the NLS estimation often produces estimates that do not satisfy the usual stability condition for each of the regimes, and thus allowing for instability (but penalizing it) facilitates utilization of the NLS estimates in our three-step procedure. Step 2 is comparable to Step 2 of Lanne et al. (2017), but the multimodality of the (penalized) log-likelihood function is taken into account by making use of a robust estimation algorithm that is able to escape from local maxima. Step 3 finalizes the estimation by initializing a standard gradient based optimization algorithm from the initial estimates obtained from the previous steps, often converging to a nearby local maximum of the (penalized) log-likelihood function.
@@
Due to the high multimodality of the (penalized) log-likelihood function, Steps 2 and 3 should be ran a large number times to improve the reliability of the results. Since multiple estimation rounds are run, the estimation procedure produces a set of estimates, presumably corresponding to different local maximums of the (penalized) log-likelihood function. If there are multiple solutions with (penalized) log-likelihoods almost equal to the greatest found (penalized) log-likelihood, the blended identification procedure discussed in Section 3.5 can be employed to obtain the best local solution that facilitates labelling the shocks of interest. The proposed three-step estimation procedure is implemented in the accompanying R package sstvars (Virolainen, 2025a), which is available via the CRAN repository.
@@
To assess the performance of the penalized maximum likelihood (PML) estimator as well as our three-step estimation procedure, we conduct a small-scale Monte Carlo study, which is discussed in detail in Appendix D. Since estimation is computationally demanding, we consider simple bivariate (structural) TVAR and LSTVAR models with $p=1$ and $M=2$. To evaluate how the PML estimator performs when the AR matrices are in the penalization region, we consider two specifications of both TVAR and LSTVAR models: one where the AR matrices are well inside the stability region and one where they lie clearly in the region where the penalization term is strictly positive.
@@
According to the results (see Tables D.2 and D.3 in Appendix D for details), estimation accuracy is slightly better when the AR matrices are inside the stability region, but even when they are in the penalization region, the estimators of the AR matrices exhibit only minor small-sample bias. Additionally, estimation accuracy appears to be slightly higher for the TVAR models compared to the LSTVAR models. While for most of the parameters there is only very slight bias, we find a notable upward small-sample bias for the scale parameter, for the degrees-of-freedom parameter whose true value is not small, and for the first entry in each impact matrix. The upward biases in the scale and not-small degrees-of-freedom parameters can both be explained by weak discrimination between large and very large estimates, as large values already imply nearly discrete regime-switches and near-Gaussian distributions, respectively. The upward bias in the first entry of each impact matrix, in turn, could be explained by its sensitivity to extreme values induced by the related particularly heavy tailed skewed $t$-distribution. Nonetheless, the bias diminishes for all of the parameters as the sample size increases. Since also the standard deviations of the estimates diminish as the sample size increases, our results are in line with consistency of the PML estimator. Overall, the performance of the proposed PML estimator and the three-step estimation procedure seems reasonable.
@@ 5 Empirical application
Our empirical application studies the macroeconomic effects of climate policy uncertainty (CPU). In the related literature, Fried et al. (2022) develop a dynamic general equilibrium model incorporating beliefs about future climate policy, and find that an increased climate policy transition risk shifts investments toward cleaner capital while reducing total investments and output. Khalil and Strobel (2023) reach similar conclusions in their dynamic stochastic general equilibrium (DSGE) model, attributing the effects to financial institutions’ aversion to uncertainty. Huang and Punzi (2024) also use a DSGE model and find that CPU shocks cause firms to delay investments, reducing output and increasing inflation, which is supported by results from their recursively identified Bayesian VAR.
We are interested in studying how the effects of the CPU shock vary depending on the level of general economic policy uncertainty (EPU). Such variation can be accommodated in our structural STVAR model by specifying a logistic transition weight function with the (lagged) level of EPU as the switching variable. This approach allows us to capture potential nonlinearities and regime-dependent effects, providing more detailed results on the macroeconomic effects of the CPU shock than linear SVAR analysis.
@@ 5.1 Data
As a measure of real economic activity, we include the log of industrial production index (IPI), which is detrended by taking first differences. For measuring the price level, we use the logs of consumer price index (CPI) and producer price index (PPI), likewise detrended by taking first differences. Finally, we measure the monetary policy stance with the effective Federal funds rate that is replaced by the Wu and Xia (2016) shadow rate for the zero-lower-bound periods (RATE). The series of the included variables are presented in Figure 1.333The CPUI and EPUI data are retrieved from https://www.policyuncertainty.com, whereas IPI, CPI, PPI, and the Federal funds rate are retrieved from the Federal reserve bank of St. Louis database and the Wu and Xia (2016) shadow rate from the Federal reserve bank of Atlanta’s website.
Figure 1: Monthly U.S. time series covering the period from 1987:4 to 2024:12. From top to bottom, the variables are the climate policy uncertainty index (divided by $10$); economic policy uncertainty index (divided by $10$); the log-differences of industrial production index, consumer price index, and producer price index (multiplied by $100$); and the effective federal funds rate replaced by the Wu and Xia (2016) shadow rate for the zero lower bound periods. The bottom panel shows the transition weights of the High EPU Regime (Regime 2) of the fitted two-regime second-order structural LSTVAR model. The shaded areas indicate the U.S. recessions defined by the NBER.
@@ 5.2 The structural LSTVAR model
In order to specify the structural STVAR model (3.2)-(3.3), the transition weights and the distributions of the structural shocks must be specified. We employ the logistic transition weights with two regimes (Anderson and Vahid, 1998) given in Equation (2.4). They facilitate smooth transitions between the regimes and enable the analysis of how the effects of the CPU shock vary depending on the level of EPU, whose first lag is specified as the switching variable. The logistic weight function also has the convenient property that is nests the threshold weight function (Tsay, 1998) as a special case when the scale parameter tends to infinity, effectively incorporating the selection between these two weight functions within estimation.
@@
We estimate the structural LSTVAR model by PML with the three-step procedure proposed in Section 4, assuming that the structural shocks follow the independent skewed $t$-distributions. The skewed $t$-distribution can flexibly capture fat tails and skewness, and as the Gaussian distribution is obtained as a special case, with $\nu_{i}=\infty$ and $\lambda_{i}=0$, the validity of the identifying Assumption 1 can be assessed based on the estimates of these parameters. To avoid overfitting, particularly in the regime accommodating fewer observations, the autoregressive order $p=2$ is selected based on Hannan-Quinn information criterion.
@@
To deal with the problem of weak identification with respect to the ordering and signs of the columns of $B_{2}$, discussed in Section 3.3, 3.4, and 3.5, we implement the following, simplistic blended identification strategy that largely avoids making restrictive assumptions about the effects of the CPU shock. We normalize the first row of $B_{1}$ positive with CPUI ordered first and assume that in both regimes, one of the shocks, deemed as the CPU shock, has greater impact effect on the CPUI than any other shock, and this impact effect has the same sign in both regimes. Then, to support our identification, we check that the recovered CPU shock aligns with several major historical events affecting CPU. In particular, we find that President Trump’s election is preceded by large CPU shocks in October 2016 and 2024 ($17$ and $26$, respectively). The largest negative recovered CPU shock ($-10$), in turn, dates to August 2022, when the Inflation Reduction Act was passed, providing support for the clean energy transition. Among all the shocks recovered from the model, only the one deemed as the CPU shock has signs consistent with the above historical events affecting CPU.
@@
The location parameter estimate is $\hat{c}=14.48$, implying roughly that Regime 1 dominates when the level of EPU is not high, while Regime 2 starts prevailing in times of high EPU. Therefore, we label Regime 1 as Low EPU Regime and Regime 2 as High EPU Regime. The scale parameter estimate is $\hat{\gamma}=16.65$, indicating that transitions between the regimes are somewhat fast (rather than very smooth), which can be observed also from the bottom panel of Figure 1, where the fitted transition weights of High EPU Regime are presented.
@@
The estimates of the degrees-of-freedom and skewness parameters are (2.33,2.10,2.64,3.23,6.54,(2.33,2.10,2.64,3.23,6.54,( 2.33 , 2.10 , 2.64 , 3.23 , 6.54 , 2.68)2.68)2.68 ) and $(0.41,0.17,0.09,0.43,0.25,-0.12)$, respectively. Clearly, none of the shocks is close to Gaussian, suggesting that Assumption 1 is satisfied. Then, we check whether our model satisfies the stationarity condition of Theorem A.1 in Appendix A by computing an upper bound for the joint spectral radius of the companion form AR matrices of the regimes. [^4] We compute the upper bound using the JSR toolbox (Jungers, 2023) in MATLAB. The obtained upper bound is $0.98$, which is strictly less than one, implying that our model satisfies Assumption A.1 (in Appendix A) and is ergodic stationary by Theorem A.1. Based on the residual diagnostics presented in Appendix E, the adequacy of our model to capture the autocorrelation structure, conditional heteroskedasticity, and distribution of the data seems, overall, reasonable.
@@ 5.3 Generalized impulse response functions
Due to the $p$-step Markov property of our structural LSTVAR model, conditioning on the previous $p$ observations, $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$, is equivalent to conditioning on $\mathcal{F}_{t-1}$. To the study the state-dependence of the effects of the CPU shock, we follow a procedure similar to Lanne and Virolainen (2025) and calculate the GIRFs conditional on a given regime dominating when the shock arrives. Specifically, for each Regime $m\in\{1,2\}$, we take all the length $p$ histories $\boldsymbol{y}_{t-1}$ from the data for which the corresponding transition weight $\alpha_{m,t}$ is greater than $0.75$, indicating that Regime $m$ is clearly dominant (there are $366$ such histories for the Low EPU Regime and $80$ for the High EPU Regime). For each such history, we take the corresponding CPU shock recovered from the data, and compute the GIRF using the Monte Carlo algorithm described in Lanne and Virolainen (2025), Appendix B. Finally, GIRFs corresponding to shocks of different sign and size are made comparable by scaling them to correspond to a $5$ point instantaneous increase of CPUI.
@@
Figure 2: Generalized impulse response functions to the identified CPU shock $h=0,1,....,36$ months ahead. From top to bottom, the responses of CPUI, EPUI, IPI, CPI, PPI, the interest rate variable, and transition weights of the High EPU Regime are depicted in each row. The responses of IPI, CPI, and PPI are accumulated to (100×100$\times$100 ×)log-levels. The left column shows the responses to conditionally on the Low EPU Regime dominating when the shock arrives, whereas the right column shows the responses conditionally on the High EPU Regime dominating. All the GIRFs have been scaled so that the instantaneous increase in CPUI is $5$ points. Each GIRF is computed with a Monte Carlo algorithm similar to the one described in Lanne and Virolainen (2025, Appendix B).
@@
The distribution of the GIRFs in each regime is illustrate in a so-called ”shotgun plot” in Figure 2, which depicts each GIRF using a level of opacity such that the darkness of a region displays the concentration of GIRFs in it (cf. Lanne and Virolainen, 2025). The response of CPUI to a positive CPU shock follows a similar pattern in both regimes in the vast majority of the GIRFs, while EPUI temporarily increases in the Low EPU Regime and decreases in the High EPU Regime. A possible explanation for the decrease in the EPUI in the High EPU Regime could be that under already heightened EPU, an unexpected increase in CPU may induce a shift in focus from broader economic policy issues to climate policy.
@@
Consistent with the previous literature (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024), a positive CPU shock decreases production in both regimes. Huang and Punzi (2024) attribute the decrease in production to firms postponing investments, whereas Fried et al. (2022) and Khalil and Strobel (2023) consider mechanisms that shift investments toward cleaner capital. Although many of the GIRFs exhibit a short-term increase of IPI in the High EPU Regime, the negative long-term response of IPI seems, on average, roughly similar in both of the regimes.
@@
In both regimes, a positive CPU shock increases consumer prices, in line with Huang and Punzi (2024) and Khalil and Strobel (2023). Huang and Punzi (2024) found that a positive CPU shock increases inflation by acting as a negative supply shock, whereas according to Khalil and Strobel (2023), a positive CPU shock may decrease aggregate demand due to risk-averse households raising savings, while firms raise their prices pre-emptively to lower the probability of being stuck with a negative markup. However, we find that the inflationary effects are stronger in the High EPU Regime than in the Low EPU Regime. One possible explanation is that the firms are more risk averse in the periods of high EPU, and thereby react more sensitively. Producer prices, in turn, mostly decrease in the Low EPU Regime but increase in the High EPU Regime.
@@
The response of the interest rate variable seems consistent with the inflationary effects of the CPU shock: it increases in both regimes, but clearly more so in the High EPU Regime. Finally, consistent with the GIRFs of EPUI, a positive CPU shock temporarily increases the transition weights of the High EPU Regime in the periods of low EPU, while the effect is opposite in the periods of high EPU. Overall, our results suggest that a positive CPU shock decreases production and increases inflation in both regimes, but its inflationary effects are stronger in the periods of high EPU than in times of low EPU.